# BARK: A Fully Bayesian Tree Kernel for Black-box Optimization

### v1

@@ Appendix C Sampling time for BARK and BART
Appendix D Hyperparameter sensitivity
D.1 Node depth prior
@@ D.2 Number of samples
Appendix E Posterior sampling
Many of the derivations are included in Kapelner & Bleich (2016). We include all of these terms here for completeness, however derivations are omitted where they are presented in the literature. A table of symbols is provided in Appendix A.
@@ Appendix E Posterior sampling
Appendix F Regret bounds and kernel approximation
@@
F.1 General regret bounds: existing methods are not applicable
@@
F.2 Asymptotic bounds: kernel approximation fails
F.2.1 The Laplace kernel as an approximation
@@
F.2.2 Dropping split independence and Poisson assumptions
@@
Appendix G Optimizer details

### v2

@@ F.1 Posterior sampling
F.2 Optimization
We provide a comparison of the time taken to fit and optimize the models in the BO loop in Fig. 11.
BARK is slower than competing methods, taking approximately 50s to fit the model, and 100s to optimize. This is due to the combination of the expensive MCMC procedure, and the large optimization formulation. This is comparable to the fitting time for BART, and the time taken to evaluate BART on a grid of $2^{14}$ points. BARK is best applied in BO settings where the objective is expensive to evaluate (e.g. taking at least several minutes, or having a large associated financial cost). Note that the PestControl and MAX (material design) benchmarks reflect examples of such black-box functions. For settings where experiments are cheap and/or function evaluations are quick, we recommend alternate methods.
Figure 11: Time taken to fit the surrogate, optimize the acquisition, and combined time for each method used in the Bayesian optimization experiments. We compare the time taken per iteration across three benchmarks. For the XGBoostMNIST benchmark, GB-RBF is optimized by enumerating every discrete combination, leading to increased optimization times in line with BARK.
Appendix G Hyperparameter sensitivity
G.1 Node depth prior
@@ G.2 Number of samples
Appendix H Posterior sampling
Many of the derivations are included in Kapelner & Bleich (2016). We include all of these terms here for completeness, however derivations are omitted where they are presented in the literature. A table of symbols is provided in Appendix C.
@@ Appendix H Posterior sampling
Appendix I Regret bounds and kernel approximation
@@
I.1 General regret bounds: existing methods are not applicable
@@
I.2 Asymptotic bounds: kernel approximation fails
I.2.1 The Laplace kernel as an approximation
@@
I.2.2 Dropping split independence and Poisson assumptions
@@
Appendix J Optimizer details